Operation Summary for June 1, 2011 - June 3, 2011 |
|
Wednesday June 1 The day shift began with the Tevatron (TEV) in access for wet engine repairs and other maintenance, with Operations monitoring the Recycler stash (336.1E10), with the Antiproton Source (Pbar) in standby, with the Main Injector (MI) in standby, with Linac and Booster in access, with NuMI off for target replacement, with the MiniBooNE and Meson FTBF T-978 experiments off due to Linac being off, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store. During the day shift, Booster personnel reported installing the NorthStar experiment target in the MI-8 line upstream of the Booster dump. Operations resumed sending beam to Linac, Booster, Pbar, and MiniBooNE at noon. Cryo system personnel performed overhauls of the TEV sectors B3 and F3 wet engines. Vacuum personnel conducted TEV A4 vacuum pump repairs and leak checks. After NuMI personnel had completed testing their horn, operators prepared the area for access. During the evening shift, Operations established beam to T-978 at 5:25 PM and then disabled their beam at 6 PM. TEV accesses finished and Operations began the TEV turn on at 7:50 PM. Operations put the TEV into HEP shot setup at 10:10 PM and established store 8787 at 11:09 PM with an average initial luminosity (AIL) of 407.1E30. Thursday June 2 The midnight shift began with Operations monitoring store 8787 (288.92E30) and the stash (180.1E10), with Pbar stacking (8.6E10), and with MiniBooNE taking beam. During the midnight shift, operators reported that a Linac quadrupole magnet power supply’s (QPS130) output dropped to zero. Operators swapped out the control cards and brought the supply back on line at 5:06 AM. Operations established beam to T-978 at 5:06 AM. During the day shift, Operations terminated store 8787 at 10:20 AM and put the TEV into shot setup, however the TEV quenched. Operators traced the cause to the STATES front end not responding, which resulted in the collimators moving into the beam. Operators put the TEV into recovery mode. At 3 PM, the Linac 750 KeV area chopper's off tube failed. Linac personnel began repairs. A Linac beam detector also failed; this took only a few minutes of downtime to replace. During the evening shift, Linac personnel completed the chopper repairs at 4:05 PM. Operations put the TEV into shot setup at 4:38 PM. A Linac RF station (LRF4) went down due to a driver anode failure. Repairs held off shot setup for about an hour. Operations established store 8790 at 6:23 PM with an AIL of 144.1E30. Operations reported short down times due a P1 transfer line corrector magnet power supply trip and MI-52 kicker trip. Friday June 3 The midnight shift began with Operations monitoring store 8790 (57.7E30) and the stash (168.9E10), with Pbar stacking (10.7E10), and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8790 at 4:18 AM, put the TEV into shot setup, and established store 8791 at 5:28 AM with an AIL of 281.26E30. Operations established beam to T-978 at 5:45 AM. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow MI personnel to conduct a ramp study. |
|
More Information Comments and Suggestions |
| last modified 06/03/2011 email Fermilab |
| Security, Privacy, Legal | |